### v3

# Measurement-driven quantum advantages in shallow circuits

Abstract
Quantum advantage schemes probe the boundary between classically simulatable and classically intractable quantum dynamics. We explore the impact of mid-circuit measurements on the computational power of quantum circuits. To this effect, we focus on quantum sampling and introduce a constant-depth measurement-driven approach for efficiently sampling from a broad class of commuting diagonal quantum circuits and associated structured phase states, previously requiring polynomial-depth unitary circuits. By interleaving mid-circuit measurements with feed-forward in randomized “fan-out staircases”, our dynamical circuits bypass Lieb–Robinson light-cone constraints, enabling global entanglement with flexible auxiliary qubit usage on bounded-degree lattices (e.g., two-dimensional grids). The generated phase states exhibit random‑matrix statistics and anti‑concentration comparable to fully random architectures. We further demonstrate measurement-driven feature maps that distinguish phases of an extended SSH model from random eigenstates in a quantum machine-learning benchmark (reservoir computing). Technologically, our results harness mid-circuit measurements to realize quantum advantages on bounded‑degree hardware with a favorable topology. Conceptually, they provide complexity‑theoretic support for quantum speedups by mid-circuit measurements.
A leading paradigm for demonstrating quantum advantage is quantum sampling: implementing a quantum evolution and sampling from its output distribution that is believed to be classically intractable under standard assumptions Hangleiter and Eisert (2023). Notable examples include boson sampling Aaronson and Arkhipov (2011); Tillmann et al. (2013); Hamilton et al. (2017); Lund et al. (2017); Deshpande et al. (2022) and random circuit sampling Arute and others (2019); Bouland et al. (2019); Shepherd and Bremner (2009); Bremner et al. (2011, 2016, 2017); Bluvstein et al. (2024); Oszmaniec et al. (2022); Haferkamp et al. (2020); Ringbauer et al. (2025); Eisert and Preskill (2025). Paradigmatically, these tasks are central to clarifying the precise fine-print of quantum advantage, but they also expose a core tension: the deep, highly-connected circuits demanded conflict with the constraints of noisy, locally-connected hardware. This tension motivates the pursuit of hardware‑efficient routes to quantum advantage.
Instantaneous-quantum-polynomial-time (IQP) circuits–commuting diagonal circuits whose output probabilities are complex-temperature partition functions—form a central model Shepherd and Bremner (2009); Bremner et al. (2011, 2016, 2017); Bluvstein et al. (2024); Paletta et al. (2024); Jozsa et al. (2024). The argument for their classical sampling hardness connects the sampling task to the #P-hard problem of estimating single output probabilities, as an efficient sampler would imply an efficient estimator (leading to an unlikely collapse of the polynomial hierarchy). This connection requires the distribution to exhibit anti-concentration, where probabilities are not dominated by a few outcomes, precluding trivial classical strategies based on guessing dominant bitstrings. Worst-to-average reductions extend the hardness from contrived instances to typical random circuits used in experiments Hangleiter and Eisert (2023). Random, long-range IQP circuits at polynomial depth satisfy these criteria, providing strong evidence for their hardness Bremner et al. (2016, 2017). That said, experimentally realizing and witnessing an IQP sampling advantage faces two key challenges: limited hardware connectivity makes the long-range interactions needed for hardness depth-expensive Maslov et al. (2024), and noise that grows with depth renders noisy IQP circuits—even with additional controlled-NOT (CX) layers—classically simulable beyond a constant depth under natural noise models Nelson et al. (2024); Rajakumar et al. .
A closely related precedent is the quench-based architecture of Bermejo-Vega et al. Bermejo-Vega et al. (2018), which realizes constant-depth hard sampling on two-dimensional nearest-neighbor lattices using product-state inputs, short-time evolution under a translation-invariant Ising Hamiltonian, and terminal fixed-basis readout—all without mid-circuit measurements. That scheme yields families of nearest-neighbor, translation-invariant 2-local constant-depth IQP circuits, with the complexity-theoretic route proceeding through a nonadaptive measurement-based quantum computing encoding of deeper logical circuits on an enlarged two-dimensional lattice. Our goal here is complementary: we ask whether mid-circuit measurements and feed-forward can compress a class of dense, long-range IQP-type diagonal circuits to constant depth on bounded-degree hardware, using auxiliary qubits that are measured out and do not appear in the final output register.
In this work, we suggest a different approach using mid-circuit measurements, which not only addresses the depth and connectivity challenges facing standard IQP implementations but is also independently interesting. Basically, in our work, we show that constant-depth quantum circuits are substantially more powerful computationally when equipped with mid-circuit measurements compared to the unitary setting. This regime of noisy intermediate scale quantum (NISQ) devices with mid-circuit measurements has been dubbed NISQ+ Quek (2024): We present a NISQ+/NISQ computational separation under plausible assumptions, highlighting the computational power of measurements along the way. To put this into context, dynamic circuits with measurements and feed-forward have shown great potential in implementing certain deep quantum circuits and preparing highly entangled states Skinner et al. (2019); Piroli et al. (2021); Lu et al. (2022); Deshpande et al. (2024); DeCross et al. (2023); Foss-Feig et al. (2023); Buhrman et al. (2024); Iqbal et al. (2024); Zi et al. (2025); Bene Watts et al. (2025); McGinley et al. (2025); Yan et al. (2025); Bäumer et al. (2025, 2024); Baumer and Woerner (2024). By conditioning operations on measurement outcomes, these circuits create non-local “shortcuts” that enable the rapid generation of complex, long-range entanglement at minimal depth Lieb and Robinson (1972); Eisert and Osborne (2006); Bravyi et al. (2006); Bäumer et al. (2024); Baumer and Woerner (2024). Inspired by these advances, we introduce a measurement-driven “fan-out staircase” architecture that implements the dense, long-range interactions required for IQP computational hardness at constant depth, even on simple two-dimensional (2D) hardware. Importantly, we show that the output distributions from these shallow circuits retain key signatures of complexity (such as anti-concentration), providing strong evidence for their classical hardness while significantly relaxing the experimental requirements of depth and connectivity.
Beyond establishing a new route to quantum sampling advantages, our measurement-driven technique provides strongly entangled feature maps for quantum machine learning. We demonstrate its effectiveness in quantum reservoir computing, where a fixed quantum evolution acts as a feature map while only a simple classical readout is trained Fujii and Nakajima (2017); Nakajima et al. (2019). Even under a realistic noise model, our measurement-based reservoir—which leverages global entanglement created at constant depth—accurately classifies distinct topological phases, surpassing conventional local reservoirs.
The upshot of the present work is as follows. Mid-circuit measurements are known to enable the preparation of symmetry-protected topological states Buhrman et al. (2024); Iqbal et al. (2024) and long-range entangled states Lu et al. (2022); Sang et al. (2021) that elude constant-depth unitary circuits. They also underpin quantum advantage in certain random shallow Clifford architectures Bene Watts et al. (2025). Extending these insights has proved difficult. Here we provide some of the clearest algorithmic-level evidence that measurements with real-time feed-forward enhance shallow quantum algorithms. We do so by compressing dense IQP circuits to constant depth and demonstrating robust anti-concentration, noise resilience, and a provable expressivity advantage in a quantum machine-learning benchmark.
Figure 1: Schematic of measurement-based fan-out staircases for dense IQP sampling. (a) Multi-loop fan-out staircases (alternating-direction ladders) generated by random dynamic circuits with measurement and feed-forward. (b) Conceptual bipartite (system–auxiliary) coupling graph. Our geometrically local, bounded-degree 2D layouts are subgraphs of this schematic. (c) Constant-depth implementation of dense IQP sampling via interleaved phase rotations and randomized fan-out staircases, where measurement outcomes inform subsequent corrections. Staircase construction and transfer-matrix feed-forward are detailed in the End Matter and SM.
Measurement-prepared Hamiltonian phase states.
The core of our protocol is a constant-depth, measurement-driven method for implementing the dense, many-body Pauli-$Z$ rotations that constitute computationally hard IQP circuits. While one could conventionally generate such a term by conjugating a single-qubit rotation with a deep unitary $\mathrm{CX}$ circuit, this requires polynomial depth on locally connected hardware.
Our measurement-driven “fan-out staircase” (FS) blocks–layered circuits of fan-out gates that build global correlations–achieve this same goal in constant depth. The mechanism of a block involves three conceptual steps. First, we locally entangle system qubits with nearby auxiliary qubits using a short pattern of nearest-neighbor CX gates. Next, we measure the auxiliaries in the $X$ basis. An $X$-basis outcome $m_{k}=1$ is equivalent to having applied a Pauli-$Z$ on that auxiliary qubit just before its measurement. The resulting Pauli-$Z$ correction string on the system qubits is determined by tracking how this “hypothetical” $Z$ operator is transformed under conjugation by the full entangling unitary. This transformation—calculated by iteratively applying commutation rules (e.g., $\mathrm{CX}(c,t)\,Z_{t}=Z_{c}Z_{t}\,\mathrm{CX}(c,t)$) through the circuit’s layers—establishes a linear map from the measurement-outcome vector $\mathbf{m}$ to the correction string, captured by a binary transfer matrix $\mathcal{T}$. The exact FS construction is characterized by a depth parameter $\mathfrak{D}$ and connectivity parameters $\mathfrak{r}_{1},\mathfrak{r}_{2}$, as detailed in the End Matter (Protocol 1) and illustrated in Fig. 1(a).
By interleaving $\mathfrak{L}$ layers of FS blocks and single-$Z$ rotations, the full circuit is
$\mathcal{C}_{\mathrm{MD}}=\bigotimes_{j=1}^{n}e^{\mathrm{i}\vartheta_{\mathfrak{L}+1,j}Z_{j}}\left(\prod_{i=1}^{\mathfrak{L}}\mathsf{FS}_{i}\;\bigotimes_{j=1}^{n}e^{\mathrm{i}\vartheta_{i,j}Z_{j}}\right),$ (1)
We absorb the Pauli‑frame corrections by classically updating the subsequent rotation angles via $\tilde{\vartheta}_{i+1,j}\coloneqq\vartheta_{i+1,j}+\frac{\pi}{2}\,(\mathcal{T}^{(i)}\mathbf{m}^{(i)})_{j}\pmod{2\pi}$, as depicted in Fig. 1(c). Since an FS block conjugates single-qubit $Z$’s into $Z$-strings, the constant-depth protocol effectively realizes an IQP circuit $\mathcal{C}_{\mathbf{A},\bm{\vartheta}}=\exp\left(\mathrm{i}\sum_{i=1}^{s}\vartheta_{i}\bigotimes_{j=1}^{n}Z^{\mathbf{A}_{i,j}}\right)$ and prepares the Hamiltonian phase state vectors $|\psi_{\mathbf{A},\bm{\vartheta}}\rangle=\mathcal{C}_{\mathbf{A},\bm{\vartheta}}|+\rangle^{\otimes n}$. These are defined by a binary matrix $\mathbf{A}\in\{0,1\}^{s\times n}$, hereafter termed the architecture matrix, which our protocol constructs row by row from the Pauli-$Z$ string generated by FS blocks, generating random IQP architectures exhibiting anti-concentration.
Proposition 1 (Anti-concentration).
Fully random architectures $\mathbf{A}\in\mathbb{Z}_{2}^{s\times n}$ achieve anti-concentration with $s\in\mathcal{O}(n)$.
Restricting the angles leads to our first main result.
Theorem 1 (Measurement-driven short IQP circuits).
Fix $k\geq 1$. For angles $\bm{\vartheta}\in\{\ell\pi/2^{k}\}_{\ell=0}^{2^{k}-1}$, the measurement-driven circuit $\mathcal{C}_{\mathrm{MD}}$ implements a dense $k$-local IQP circuit.
Here, a $k$-local IQP circuit consists of commuting $Z$-interactions involving at most $k$ qubits (i.e., weight $\leq k$). We call it dense if an $n$-independent constant fraction of admissible $Z$-strings appear with $\Theta(1)$ angles. In particular, the diagonal gates $e^{\mathrm{i}\pi Z/8}$ and $e^{\mathrm{i}\pi ZZ/4}$ arise from the synthesis with $k=3$ and $k=2$, respectively, using a constant number of FS blocks on bounded-degree 2D layouts. To quantitatively evaluate the statistical randomness of binary circuit architectures, we propose the following criterion:
Figure 2: (a) Checkerboard layout of alternating system and auxiliary qubits on a 2D square lattice, with a sampled Hamiltonian path covering all qubits. (b) Comparison of eigenvalue distributions from randomized fan-out staircases with the theoretical Marchenko–Pastur law for a representative $41\times 41$-qubit system. (c, d) Minimal circuit depths required to satisfy Criterion 1, contrasting measurement-driven fan-out staircases (blue squares) with circuits without auxiliary qubits (red circles), for (c) 2D grid and (d) all-to-all connectivity.
Criterion 1: $\mathbf{A}\in\mathbb{Z}_{2}^{s\times n}$ is statistically random if (i) the eigenvalue distribution of the standardized covariance matrix $(2\mathbf{A}-1)(2\mathbf{A}-1)^{\top}/s$ matches Marchenko–Pastur law Marchenko and Pastur (1967); (ii) row and column Hamming weights and pairwise distances follow binomial distributions; and (iii) $\geq 90\%$ of random submatrices exhibit ranks within 2 of full rank over GF(2).
Remark. Our randomness diagnostics probe typicality only—analogous to statistical validation in boson and random-circuit sampling—and do not imply a worst-to-average reduction Hangleiter and Eisert (2023). Such a reduction would require distribution-specific random self-reducibility, which we leave open.
These conditions guarantee spectral, statistical, and algebraic universality Marchenko and Pastur (1967); Edelman and Rao (2005); Tao (2012). Using Criterion 1, we numerically compare measurement-driven randomized FS against random $\mathrm{CX}$ circuits without auxiliary qubits under both 2D grid and all-to-all connectivity. In the 2D configuration, system and auxiliary qubits alternate, with random directed Hamiltonian paths generated according to Protocol 2 in the End Matter, respecting local constraints [Fig. 2(a)]. Fig. 2(b) shows eigenvalue distributions from measurement-driven FS closely follow the Marchenko–Pastur law, unlike random local $\mathrm{CX}$ circuits at comparable depths. Numerical results [Fig. 2(c,d)] indicate that circuits without auxiliary qubits require polynomial (2D grid) or logarithmic (all-to-all) depths to satisfy Criterion 1. In contrast, measurement-driven FS meet this criterion at constant depth, with parameters $\mathfrak{r}_{1}=\mathfrak{r}_{2}=1$, and $\mathfrak{D}=2$ for all-to-all connectivity and $\mathfrak{D}=4$ for 2D grid connectivity, even in the large-scale regime. Subsequent IQP simulations employ the 2D grid connectivity, interleaving single-qubit $Z$ rotations with two FS layers (each with $\mathfrak{D}=2$).
Quantum circuit cost quantifies the minimal resources required to implement a given unitary operation by integrating contributions from local gate interactions, where the entanglement entropy generated across suitable 1D bipartitions provides a fundamental lower bound for this cost Mariën et al. (2016); Eisert (2021). We extend this to 2D lattices by considering horizontal $(\ell_{x})$ and vertical $(\ell_{y})$ cuts, yielding the summed entanglement entropy
$\xi(\psi):=\frac{1}{\eta}\biggl(\sum_{\ell_{x}}S(\psi,\ell_{x})+\sum_{\ell_{y}}S(\psi,\ell_{y})\biggr),$ (2)
where $\eta$ is a geometry-dependent constant. $\xi$ provides a lower bound for the geometrically local circuit cost required to prepare a given state from a product in two dimensions (see the SM). We benchmark against $\xi_{\mathrm{lin}}$, the average for output states of deep, linear-depth random CX circuits ($\sim 6n$). Although such constant-depth circuits on 2D lattices are not expected to achieve genuine volume-law entropy across each cut—particularly for larger systems—Fig. 3(a) shows that measurement-driven IQP circuits nonetheless achieve $\xi/\xi_{\mathrm{lin}}\approx 1$, matching the entangling power of linear-depth CX circuits. In contrast, conventional depth-matched IQP circuits without auxiliary systems fall significantly short.
The collision probability, $\chi:={\mathbb{E}}_{\mathcal{C}}\left[\sum_{x}p_{\mathcal{C}}(x)^{2}\right]$, quantifies quantum circuit anti-concentration. For sufficiently deep Haar-random circuits, the collision probability approaches $\chi_{\text{Haar}}=2/(2^{n}+1)$. Anti-concentration requires the ratio $\chi/\chi_{\text{Haar}}$ to remain bounded by a small constant Dalzell et al. (2022), a condition underpinning complexity arguments and benchmarks like cross-entropy benchmarking Boixo et al. (2018); Arute and others (2019). Fig. 3(b) compares collision probabilities between measurement-driven IQP circuits and standard IQP circuits without auxiliary qubits at comparable depths. Unlike standard IQP circuits, measurement-driven IQP circuits exhibit a decreasing collision probability ratio with increasing system size, indicating robust anti-concentration at constant depth and minimal sensitivity to connectivity constraints.
Building upon (i) Prop. 1, (ii) our 2D numerics, and (iii) Lemma 8 of Ref. Hangleiter et al. (2018), which establishes #P-hardness for depth-$\mathcal{O}(n)$ random circuits over $\{e^{\mathrm{i}\pi Z/8},\,e^{\mathrm{i}\pi ZZ/4},\mathrm{SWAP}\}$, we make the following observation. On bounded-degree 2D layouts with an interleaved system–auxiliary placement, our constant two-qubit depth protocol implements a subclass of dense $k$-local diagonal layers, including the gates $e^{\mathrm{i}\pi Z/8}$ and $e^{\mathrm{i}\pi ZZ/4}$ central to that hardness result (Thm. 1). Since the SWAP permutations also required by the full construction in Ref. Hangleiter et al. (2018) are realizable in our model via measurement-mediated routing without added unitary depth, our scheme therefore realizes a subset of instances from this known-hard circuit family at constant depth. This serves as strong motivation—not a formal worst-to-average reduction Bouland et al. (2019); Haferkamp et al. (2020); Bremner et al. (2016)—for the more permissive conjecture that follows.
Figure 3: Comparison of measurement-driven IQP sampling (blue) with standard IQP sampling (red/gray) at matched constant two-qubit depth on a 2D nearest-neighbor grid. (a) Relative entanglement-based circuit cost $\xi/\xi_{\mathrm{lin}}$ for fixed-depth circuits $\mathfrak{L}=2,\mathfrak{D}=2$ versus system size. (b) Collision probability ratio $\chi/\chi_{\text{Haar}}$ versus system size. (c) Total variation distance between noisy and ideal output distributions under gate depolarizing errors. (d) Total variation distance under pure dephasing noise assuming fixed gate duration.
Conjecture 1 (Average-case hardness).
For a broad class of connected bounded‑degree bipartite lattices in dimension $d\geq 2$, there exists a constant‑two‑qubit‑depth, measurement‑driven IQP family—realized via randomized fan‑out staircases on $n$ qubits with at most $O(n\log n)$ auxiliaries and i.i.d. phases in $0,2\pi)$—whose output distributions anti‑concentrate and for which, on a non‑negligible fraction of instances, at least one fixed output probability is #P‑hard to approximate within a constant multiplicative factor.
Remark. Our numerics suggest $\mathcal{O}(n)$ auxiliaries often suffice on 2D grids; the $\mathcal{O}(n\log n)$ budget in the conjecture is a conservative high‑probability allowance to guarantee strong pseudo‑randomness across a wider range of bounded‑degree layouts (e.g., by modestly widening fan‑out targets or layering additional auxiliary resources), at constant two‑qubit depth.
Implication. Assuming Conjecture 1 and non-collapse of the polynomial hierarchy, the standard Stockmeyer route implies that approximate sampling from this family is classically intractable up to inverse-polynomial total-variation error Hangleiter and Eisert (2023); Stockmeyer (1983).
Measurement-driven IQP circuits exhibit notably enhanced noise resilience compared to conventional IQP circuits, which typically become classically simulable under constant gate noise beyond certain depth thresholds Nelson et al. (2024); Rajakumar et al. . By compressing a family of polynomial-depth circuits into constant depth, measurement-driven circuits evade these simulation methods typically applicable to deeper noisy circuits.
To quantify this advantage, we numerically evaluate the total variation distance $\delta_{\mathrm{TV}}(p,\widetilde{p}):=\sum_{x}\left|p(x)-\widetilde{p}(x)\right|/2$ between the ideal output distribution $p(x)$ and the noisy distribution $\widetilde{p}(x)$ for standard and measurement-driven IQP circuits on a $6\times 6$ 2D grid, assuming ideal measurements. We first analyze gate depolarizing noise at varying error rates (Fig. 3(c)). Measurement-driven circuits consistently yield lower $\delta_{\mathrm{TV}}$ compared to standard IQP circuits, primarily due to efficient long-range interactions without additional $\mathrm{SWAP}$ gates. Moreover, measurement-driven circuits exhibit narrower $\delta_{\mathrm{TV}}$ distributions, as local errors propagate into correlated global errors through feed-forward corrections.
In a second analysis, we omit gate errors and focus exclusively on decoherence characterized by the dephasing coherence time $T_{2}$ (assuming an infinite $T_{1}$). The total variation distance $\delta_{\mathrm{TV}}$ is expected to saturate exponentially with circuit duration $\Delta t$ according to: $\delta_{\mathrm{TV}}(\Delta t)\approx\delta_{\mathrm{TV},\infty}(1-e^{-\kappa\Delta t})$, where $\delta_{\mathrm{TV},\infty}$ is the asymptotic distance between the ideal and uniform distributions, and $\kappa$ is an effective dephasing rate. Numerical results (assuming 200 ns per $\mathrm{CX}$ layer) in Fig. 3(d) show that depth compression in measurement-driven circuits significantly reduces $\delta_{\mathrm{TV}}$, demonstrating substantial mitigation of decoherence effects.
Measurement-based quantum reservoir.
Reservoir computing (RC) embeds inputs using a fixed dynamical system while training only a simple classical readout Jaeger (2001); Maass et al. (2002). In quantum RC, the reservoir is a fixed quantum evolution whose intrinsic dynamics generate rich features, with measurements inducing nonlinearity and entanglement enabling long-range mixing Fujii and Nakajima (2017); Nakajima et al. (2019). In our scheme, interleaving single-qubit gates with $\mathsf{FS}$ and $\mathsf{FS}^{\dagger}$ forms a constant-depth reservoir where mid-circuit measurements and Pauli-frame feed-forward create input-dependent nonlinear feature maps.
We benchmark our measurement-based quantum reservoir to an extended bosonic Su–Schrieffer–Heeger chain, $\mathcal{H}=\sum_{\mu\in\{x,y,z\}}\gamma_{\mu}\left(J\sum_{i=1}^{n/2}\sigma_{2i-1}^{\mu}\sigma_{2i}^{\mu}+J^{\prime}\sum_{i=1}^{n/2-1}\sigma_{2i}^{\mu}\sigma_{2i+1}^{\mu}\right)$, where $\gamma_{\mu}=1+(\delta-1)\delta_{\mu,z}$. This model features three distinct phases—trivial, topological, and symmetry-broken—depending on the values of $J$, $J^{\prime}$, and $\delta$ Su et al. (1979); de Léséleuc et al. (2019); Elben et al. (2020); Cao et al. (2025). For each phase, we sample 500 eigenstates uniformly from the lowest 20 levels, apply weak local perturbations, and evolve under the candidate reservoirs. Only local spin observables $\langle\sigma^{z}_{i}\rangle$ are measured for each reservoir (8192 shots), with a simulated symmetric per‑qubit readout error of $0.5\%$.
We compare four reservoirs: (i) a random anisotropic Heisenberg reservoir, $\mathcal{H}_{\mathrm{H}}=\sum_{\langle i,j\rangle}\sum_{\mu}J_{i,j}^{\mu}\,\sigma^{\mu}_{i}\sigma^{\mu}_{j}$, where $\mu\in\{x,y,z\}$ labels spin components; (ii) a random transverse-field Ising reservoir, $\mathcal{H}_{\mathrm{TFI}}=\sum_{\langle i,j\rangle}J_{i,j}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{x}_{i}$; (iii) a random transverse-field XY reservoir $\mathcal{H}_{\mathrm{XY}}=\sum_{\langle i,j\rangle}(J^{x}_{i,j}\sigma^{x}_{i}\sigma^{x}_{j}+J^{y}_{i,j}\sigma^{y}_{i}\sigma^{y}_{j})+\sum_{i}h_{i}\sigma^{x}_{i}$; and (iv) a measurement-based multibody XY reservoir $\mathcal{H}_{\mathbf{A}}=\sum_{i=1}^{s}(c^{x}_{i}\prod_{j\in A_{i}}\sigma^{x}_{j}+c^{y}_{i}\prod_{j\in A_{i}}\sigma^{y}_{j})$, with all coupling coefficients $J_{i,j}^{x,y,z}$, $h_{i}$, and $c_{i}^{x,y,z}$ drawn i.i.d. from $\mathcal{N}(0,1/2)$, $s=16$, and $\mathbf{A}$ encodes the measurement‑enabled architecture. All reservoirs use 16 system qubits on 2D-square and hexagonal lattices, undergo ten identical Floquet cycles generated by their Hamiltonian terms. The classification of phases is performed using a support vector machine (SVM) with a radial basis function kernel Cortes and Vapnik (1995); Chang and Lin (2011).
Figure 4: Phase-classification accuracy versus Floquet cycle for different reservoirs. Each violin distribution is obtained from 50 random realizations of architectures and coupling strengths. The measurement-driven multibody XY reservoir achieves high accuracy at short times, outperforming local reservoirs (i)-(iii), while removing feed-forward (FF) adaptivity significantly reduces performance.
The classification accuracy versus Floquet cycle for the reservoirs is shown in Fig. 4, with each violin distribution obtained from 50 independent random realizations of $\mathbf{A}$ and coupling strengths. In contrast to local reservoirs (i)–(iii), the measurement‑driven multibody reservoir rapidly reaches high accuracy at short times. However, removing the feed‑forward (FF) adaptivity leads to a marked drop in performance. From an expressivity standpoint, local-Hamiltonian reservoirs with constant Floquet cycles cannot approximate certain measurement-based feature maps even within inverse-polynomial error, a distinction formalized in Thm. 2 (proof and additional benchmarks in the SM).
Theorem 2 (Expressivity separation for local Hamiltonian reservoirs).
Let $G$ be a connected bounded‑degree graph with $n$ qubits and fix a Floquet period $T$. There exist encoded input state vectors $|\psi_{0}\rangle,|\psi_{1}\rangle$, a two‑outcome observable $O$, and a measurement‑driven Floquet reservoir $\mathcal{R}_{\mathrm{MD}}^{(T)}$ such that
[…]
while any nearest-neighbor Hamiltonian-based Floquet evolution on the same graph $G$, with the same total evolution time, achieves a separation at most exponentially small in $n$.
Discussion and outlook.
Our measurement-driven framework incorporates extensive interactions between system and auxiliary qubits to broaden the scope of dynamic quantum circuits. In contrast to depth-heavy random-unitary experiments Arute and others (2019), our dynamic-circuit scheme achieves anti-concentration at constant two-qubit depth via mid-circuit measurements with modest auxiliary overhead. Beyond sampling, its associated Hamiltonian-phase-state preparations can enhance cryptographic protocols Bostanci et al. (2024). An extension to fermion sampling is also promising, as mid-circuit measurements can introduce the non-Gaussian effects crucial for computational hardness Oszmaniec et al. (2022). Furthermore, our measurement-based feature maps can advance quantum machine learning applications by enabling constant-depth global operations that capture data structures inaccessible to local unitary kernels Schuld and Killoran (2019); Havlíček et al. (2019); Gil-Fuster et al. (2024).
[…]
Regarding quantum error correction, our shallow-circuit techniques offer a promising route to enhance quantum coding protocols based on low-depth random circuits Gullans et al. (2021). Our approach generalizes dynamic-circuit methods for preparing toric code ground states with long-range topological order Verresen et al. (2021); Foss-Feig et al. (2023), and may enable efficient realization of broader classes of quantum low-density parity-check (LDPC) codes Breuckmann and Eberhardt (2021), thereby providing a viable blueprint for scalable and fault-tolerant quantum-information processing.
Conclusion.
We introduced an explicit, measurement-based scheme achieving dense IQP sampling at constant depth via randomly generated fan-out staircases, bypassing conventional light-cone constraints to enable global entanglement, robust anti-concentration, and high noise resilience. Additionally, we demonstrated that measurement-based circuits effectively serve as quantum reservoirs, efficiently distinguishing topological phases from random eigenstates. Overall, our results highlight that mid-circuit measurements and feed-forward can significantly enhance quantum algorithm performance at shallow depth—alleviating connectivity constraints and opening pathways for NISQ+ quantum advantages as next steps to be taken towards achieving fault-tolerant application-scale Eisert and Preskill (2025) quantum computers.
Acknowledgments.
Acknowledgements.
We thank Antonio Anna Mele, Marcel Hinsche, Elies Gil-Fuster, Zhenhuan Liu, Julio Carlos Magdalena De La Fuente, and Xiaoyu He for insightful discussions and valuable feedback. This work has been supported by the BMFTR (DAQC, MuniQC-Atoms), Clusters of Excellence (ML4Q, MATH+), the Munich Quantum Valley, Berlin Quantum, the Quantum Flagship (MILLENION, PASQUANS2), the DFG (CRC 183, SPP 2514), the European Research Council (DebuQC), and the Alexander-von-Humboldt Foundation.
Supplementary Material for “Measurement-driven quantum advantages in shallow circuits”
Appendix A Appendix A: Fan-out ladders and staircases
Figure S1: (a) Schematic of a one-dimensional topology where system qubits and auxiliary qubits alternate. (b) A $\mathrm{CX}$ ladder, denoted $\mathsf{FL}_{1}$, implemented in a 1D topology with one layer. (c–g) Fan-out ladders denoted $\mathsf{FL}^{(\uparrow)}_{(1,2)}$$\mathsf{FL}^{(\uparrow)}_{(2,3)}$$\mathsf{FL}^{(\uparrow)}_{(2,3,4)}$$\mathsf{FL}^{(\uparrow)}_{(1,2,4,5)}$, and $\mathsf{FL}^{(\uparrow)}_{(1,4,5,6)}$ are achieved through 2, 3, 4, 5, and 6 repetitions of nearest-neighbor $\mathrm{CX}$ layers, respectively. In these 1D implementations, each successive fan-out operation does not necessarily target a contiguous block of qubits—some intermediate connections are partially canceled. Nevertheless, for a fixed number of layers and away from circuit boundaries, the overall structure retains translational symmetry.
Figure S2: (a) A bipartite graph topology in which each system qubit interacts with multiple nearest‐neighbor auxiliary qubits. (b) A dynamic circuit realizing the fan‐out ladder $\mathsf{FL}^{(\uparrow)}_{(1,2,3,4)}$, where each system qubit $Q_{i}$ is connected to auxiliary qubits $\widetilde{Q}_{i-2}$, $\widetilde{Q}_{i-1}$, $\widetilde{Q}_{i}$, and $\widetilde{Q}_{i+1}$ (when available). (c) A dynamic circuit realizing the fan‐out ladder $\mathsf{FL}^{(\uparrow)}_{(1,4,7,8)}$, where each system qubit $Q_{i}$ is connected to auxiliary qubits $\widetilde{Q}_{i-4}$, $\widetilde{Q}_{i-1}$, $\widetilde{Q}_{i}$, and $\widetilde{Q}_{i+3}$ (when available). (d) A dynamic circuit realizing the fan‐out ladder $\mathsf{FL}^{(\uparrow)}_{(3,5,7,10)}$, where each system qubit $Q_{i}$ is connected to auxiliary qubits $\widetilde{Q}_{i-4}$, $\widetilde{Q}_{i-2}$, $\widetilde{Q}_{i+1}$, and $\widetilde{Q}_{i+3}$ (when available). (e) A dynamic circuit realizing the fan‐out ladder $\mathsf{FL}^{(\uparrow)}_{(1,2,3,5,\ldots,12)}$, where each system qubit $Q_{i}$ is connected to auxiliary qubits $\widetilde{Q}_{i-4}$, $\widetilde{Q}_{i-2}$, $\widetilde{Q}_{i-1}$, $\widetilde{Q}_{i}$, $\widetilde{Q}_{i+1}$, and $\widetilde{Q}_{i+3}$ (when available). In all cases, the structure of each fan‐out gate is determined by the propagation of information in the original circuit, and the overall configuration retains translational symmetry away from circuit boundaries.
Figure S3: (a) A two-layer grid topology with alternating system and auxiliary qubits, containing $2k$ system qubits and 2k12k-1 auxiliary qubits for an integer $k$. (b) A three-layer grid topology with alternating system and auxiliary qubits, containing 3k13k-1 system qubits and 3k13k-1 auxiliary qubits for an integer $k$. (c) A $2k\times 2k$ grid topology (with one missing corner) containing $2k^{2}$ system qubits and $2k^{2}-1$ auxiliary qubits. (d) An example 15-qubit dynamic circuit based on the grid topology in (a) that realizes a non-translationally invariant (irregular) fan-out ladder. (e) An example 15-qubit dynamic circuit based on the grid topology in (b) that realizes an irregular fan-out ladder. (f) An example 15-qubit dynamic circuit based on the grid topology in (c) that realizes an irregular fan-out ladder.
Fan-out staircases are constructed by concatenating multiple fan-out ladders, with adjacent ladders arranged in opposite orders. In this appendix, we describe in detail how to construct fan-out ladders that implement various fan-out gates under different connectivity topologies. We also explain how these fan-out ladders can be combined to form a fan-out staircase that incorporates multiple rounds of mid-circuit measurement while requiring only a single round of feed-forward. Figs. S1, S2, and S3 illustrate some fan-out ladder examples in one-dimensional (1D) connectivity, bipartite graph connectivity, and two-dimensional (2D) grid topologies, respectively.
A.0.1 One-dimensional topology
Consider first the 1D connectivity case, where system qubits and auxiliary qubits alternate, as shown in Fig. S1(a). It is well known that a single layer of nearest-neighbor $\mathrm{CX}$ operations, combined with mid-circuit measurements, can implement a $\mathrm{CX}$ ladder [Fig. S1(b)]. (Apparently, the inverse of this ladder can be used to prepare GHZ states.) By repeating these operations multiple times, one obtains fan-out ladders that—despite partial cancellation of some intermediate connections—retain overall translational symmetry away from circuit boundaries. The detailed structure of the required feed-forward operations is determined by the circuit connectivity. In what follows, we rigorously derive the equivalent transformation implemented by the $\mathrm{CX}$/fan-out ladder and characterize the corresponding feed-forward corrections based on the measurement outcomes for one and two repetitions of the $\mathrm{CX}$ layers.
Single repetition of the $\mathrm{CX}$ layer.
Let the system qubits be denoted by
$Q_{1},\,Q_{2},\,\dots,\,Q_{n},$ (S1)
and the auxiliary qubits by
$\widetilde{Q}_{1},\,\widetilde{Q}_{2},\,\dots,\,\widetilde{Q}_{n-1},$ (S2)
with each auxiliary qubit initially prepared in the state vector $|0\rangle$. We denote the controlled-NOT (CX) gate with control qubit $Q_{i}$ and target qubit $\widetilde{Q}_{j}$ as $\mathrm{CX}(Q_{i},\widetilde{Q}_{j})$, and the reverse operation, where $\widetilde{Q}_{j}$ is the control and $Q_{i}$ is the target, as $\mathrm{CX}(\widetilde{Q}_{j},Q_{i})$.
The overall unitary corresponding to a single repetition of the $\mathrm{CX}$ layer is given by
$U^{(1)}=\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)\cdot\left(\prod_{i=1}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right).$ (S3)
We now show that if, for a given auxiliary qubit $\widetilde{Q}_{i}$ (initialized in $|0\rangle$), we sequentially apply $\mathrm{CX}(Q_{i},\widetilde{Q}_{i})$ and $\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})$, then subsequently applying a Hadamard gate on $\widetilde{Q}_{i}$ and measuring $\widetilde{Q}_{i}$ in the computational basis with outcome $|0\rangle$ yields an effective operation on the system qubits equivalent to a $\mathrm{CX}$ gate from $Q_{i}$ to $Q_{i+1}$. For clarity, we now detail the state evolution.
Assume that the initial state vector of the three qubits involved is
$|\Psi_{0}\rangle=\Bigl(\alpha\,|0\rangle_{Q_{i}}+\beta\,|1\rangle_{Q_{i}}\Bigr)\otimes|0\rangle_{\widetilde{Q}_{i}}\otimes\Bigl(\gamma\,|0\rangle_{Q_{i+1}}+\delta\,|1\rangle_{Q_{i+1}}\Bigr),$ (S4)
where $\alpha,\beta,\gamma,\delta\in\mathbb{C}$ satisfy the normalization conditions.
The application of the first gate, $\mathrm{CX}(Q_{i},\widetilde{Q}_{i})$, transforms the state vector into
$|\Psi_{1}\rangle={}$ $\alpha\,|0\rangle_{Q_{i}}\,|0\rangle_{\widetilde{Q}_{i}}\,\Bigl(\gamma\,|0\rangle_{Q_{i+1}}+\delta\,|1\rangle_{Q_{i+1}}\Bigr)$ (S5)
$+\beta\,|1\rangle_{Q_{i}}\,|1\rangle_{\widetilde{Q}_{i}}\,\Bigl(\gamma\,|0\rangle_{Q_{i+1}}+\delta\,|1\rangle_{Q_{i+1}}\Bigr).$
Subsequently, the application of the second gate, $\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})$, results in the state vector
$|\Psi_{2}\rangle={}$ $\alpha\,|0\rangle_{Q_{i}}\,|0\rangle_{\widetilde{Q}_{i}}\,\Bigl(\gamma\,|0\rangle_{Q_{i+1}}+\delta\,|1\rangle_{Q_{i+1}}\Bigr)$ (S6)
$+\beta\,|1\rangle_{Q_{i}}\,|1\rangle_{\widetilde{Q}_{i}}\,\Bigl(\gamma\,|1\rangle_{Q_{i+1}}+\delta\,|0\rangle_{Q_{i+1}}\Bigr).$
We now apply the Hadamard gate on $\widetilde{Q}_{i}$ and project onto the subspace where the auxiliary qubit is measured in the computational basis with outcome $|0\rangle$. Discarding the normalization factor, the resulting (post-measurement) state is
$|\Psi_{f}\rangle\propto{}$ $\alpha\,|0\rangle_{Q_{i}}\,\Bigl(\gamma\,|0\rangle_{Q_{i+1}}+\delta\,|1\rangle_{Q_{i+1}}\Bigr)$ (S7)
$+\beta\,|1\rangle_{Q_{i}}\,\Bigl(\gamma\,|1\rangle_{Q_{i+1}}+\delta\,|0\rangle_{Q_{i+1}}\Bigr).$
This final state is equivalent to the state obtained by directly applying a $\mathrm{CX}$ gate with $Q_{i}$ as control and $Q_{i+1}$ as target to the initial two-qubit state. If all auxiliary measurement outcomes are $|0\rangle$, the overall effect is a $\mathrm{CX}$ ladder, as illustrated in Fig. S1(b).
Now we consider the more general case in which, after applying $U^{(1)}$ and the Hadamard, the measurement outcome on a particular auxiliary qubit $\widetilde{Q}_{j}$ is $|1\rangle$. This outcome is equivalent to having first applied a $Z$ gate to $\widetilde{Q}_{j}$ prior to the Hadamard transformation, followed by measurement, which then yields the outcome $|0\rangle$. Using the propagation relations for Pauli-$Z$ operators under conjugation by $\mathrm{CX}$ gates gives
$Z(\widetilde{Q}_{j})\left(\prod_{i}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right)$ (S8)
$=$ $\left(\prod_{i}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right)Z(Q_{j})Z(\widetilde{Q}_{j}),$
and
$Z(\widetilde{Q}_{j-1})Z(Q_{j})\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)$ (S9)
$=$ $\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)Z(Q_{j}),$
one obtains, by successive application,
$Z(\widetilde{Q}_{j})U^{(1)}=$ $U^{(1)}Z(Q_{j})Z(\widetilde{Q}_{j})$ (S10)
$=$ $Z(\widetilde{Q}_{j-1})Z(Q_{j})U^{(1)}Z(\widetilde{Q}_{j})$
$=$ $Z(\widetilde{Q}_{j-2})Z(Q_{j-1})Z(Q_{j})U^{(1)}Z(\widetilde{Q}_{j-1})Z(\widetilde{Q}_{j})$
$=$ $\cdots$
$=$ $\left(\prod_{i=1}^{j}Z(Q_{i})\right)U^{(1)}\left(\prod_{i=1}^{j}Z(\widetilde{Q}_{i})\right).$
Since the initial state vectors of the auxiliary qubits are $|0\rangle$, the application of $\left(\prod_{i=1}^{j}Z(\widetilde{Q}_{i})\right)$ does not alter their state. Consequently, obtaining a measurement outcome $m_{j}=1$ on $\widetilde{Q}_{j}$ is equivalent to the case $m_{j}=0$ up to the application of corrective $Z$ operations on the system qubits $Q_{1},Q_{2},\dots,Q_{j}$.
More generally, we can encapsulate the required feed-forward correction in terms of a binary transfer matrix $\mathcal{T}$ that is determined by the circuit structure. In particular, for each system qubit $Q_{i}$, the corrective $Z$ operation can be written as
$Z^{f_{i}}\quad\text{with}\quad f_{i}=\bigoplus_{j}\mathcal{T}_{i,j}\,m_{j},$ (S11)
where $m_{j}$ is the measurement outcome (0 or 1) for the auxiliary qubit $\widetilde{Q}_{j}$, and the binary matrix $\mathcal{T}$ encodes the propagation of the $Z$ operators through the circuit. In other words, the feed-forward correction on the system qubits is completely determined by the parity of the measured outcomes on the corresponding auxiliary qubits, as dictated by the transfer matrix $\mathcal{T}$. For the one-layer case, the transfer matrix $\mathcal{T}$ is given by:
$\mathcal{T}_{i,j}=\begin{cases}1,&\text{ if }i\leq j,\\ 0,&\text{ if }i>j,\end{cases}$ (S12)
where $i$ and $j$ denote the indices of the system qubits and auxiliary qubits, respectively.
Two repetitions of the $\mathrm{CX}$ layer.
We now extend the analysis to the case of two repetitions of the nearest-neighbor $\mathrm{CX}$ layers. The overall unitary for two repetitions is given by
$U^{(2)}=U^{(1)}\cdot U^{(1)}.$ (S13)
To illustrate the construction, we focus on a block comprising five qubits: system qubits $Q_{i}$, $Q_{i+1}$, $Q_{i+2}$, and auxiliary qubits $\widetilde{Q}_{i}$, $\widetilde{Q}_{i+1}$. The sequence of $\mathrm{CX}$ gates acting on these qubits is
$\mathrm{CX}(Q_{i},\widetilde{Q}_{i}),\quad\mathrm{CX}(Q_{i+1},\widetilde{Q}_{i+1}),$ (S14)
$\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1}),\quad\mathrm{CX}(\widetilde{Q}_{i+1},Q_{i+2}),$
$\mathrm{CX}(Q_{i+1},\widetilde{Q}_{i+1}),\quad\mathrm{CX}(\widetilde{Q}_{i+1},Q_{i+2}).$
After these $\mathrm{CX}$ operations, Hadamard gates are applied to both $\widetilde{Q}_{i}$ and $\widetilde{Q}_{i+1}$ followed by measurement in the computational basis.
For clarity, let the initial state vector of these five qubits be
$|\Psi_{0}\rangle=$ $|\psi\rangle_{Q_{i}}\otimes|0\rangle_{\widetilde{Q}_{i}}\otimes|\phi\rangle_{Q_{i+1}}\otimes|0\rangle_{\widetilde{Q}_{i+1}}\otimes|\chi\rangle_{Q_{i+2}},$ (S15)
with $|\psi\rangle_{Q_{i}}=\alpha\,|0\rangle+\beta\,|1\rangle$, $|\phi\rangle_{Q_{i+1}}=\gamma|0\rangle+\delta\,|1\rangle$, $|\chi\rangle_{Q_{i+2}}=\gamma^{\prime}|0\rangle+\delta^{\prime}\,|1\rangle$. where $\alpha,\beta,\gamma,\delta,\gamma^{\prime},\delta^{\prime}\in\mathbb{C}$ and the states are normalized.
Assuming that the Hadamard measurements on both auxiliary systems yield $|0\rangle$ (i.e., $m_{i}=0$ and $m_{i+1}=0$), the post-selected state vector becomes
$|\Psi_{f}\rangle\propto{}$ $\alpha\gamma\,|0\rangle_{Q_{i}}\,|0\rangle_{Q_{i+1}}\,\Bigl(\gamma^{\prime}\,|0\rangle_{Q_{i+2}}+\delta^{\prime}\,|1\rangle_{Q_{i+2}}\Bigr)$ (S16)
$+\alpha\delta\,|0\rangle_{Q_{i}}\,|1\rangle_{Q_{i+1}}\,\Bigl(\delta^{\prime}\,|0\rangle_{Q_{i+2}}+\gamma^{\prime}\,|1\rangle_{Q_{i+2}}\Bigr)$
$+\beta\delta\,|1\rangle_{Q_{i}}\,|0\rangle_{Q_{i+1}}\,\Bigl(\gamma^{\prime}\,|0\rangle_{Q_{i+2}}+\delta^{\prime}\,|1\rangle_{Q_{i+2}}\Bigr)$
$+\beta\gamma\,|1\rangle_{Q_{i}}\,|1\rangle_{Q_{i+1}}\,\Bigl(\delta^{\prime}\,|0\rangle_{Q_{i+2}}+\gamma^{\prime}\,|1\rangle_{Q_{i+2}}\Bigr).$
The net effect of these operations is equivalent to first applying a $\mathrm{CX}$ gate from $Q_{i+1}$ to $Q_{i+2}$, followed by a fan-out gate with control $Q_{i}$ and target qubits $Q_{i+1}$ and $Q_{i+2}$. By incorporating additional $\mathrm{CX}$ layers over a larger set of qubits, we obtain a deeper fan-out ladder. In the two-layer case, if all measurement outcomes are $|0\rangle$, the resulting fan-out ladder—denoted $\mathsf{FL}_{1,2}$—is shown in Fig. S1(c).
A similar analysis applies to the feed-forward operations when the measurement outcome on a particular auxiliary qubit $\widetilde{Q}_{j}$ is $|1\rangle$. As before, this outcome is equivalent to having first applied a $Z$ gate to $\widetilde{Q}_{j}$ before the Hadamard transformation, thereby converting the outcome to an effective $|0\rangle$ measurement while introducing extra $Z$ operators on the system qubits.
Using the propagation relations for Pauli-$Z$ operators under conjugation by $\mathrm{CX}$ gates, i.e., Eqs. (S8), (S9), we have
$Z(\widetilde{Q}_{j})U^{(2)}=$ $U^{(1)}Z(Q_{j})Z(\widetilde{Q}_{j})U^{(1)}$ (S17)
$=$ $U^{(1)}Z(Q_{j})U^{(1)}Z(Q_{j})Z(\widetilde{Q}_{j})$
$=$ $U^{(1)}Z(\widetilde{Q}_{j-1})U^{(1)}Z(\widetilde{Q}_{j})$
$=$ $U^{(2)}Z(\widetilde{Q}_{j-1})Z(Q_{j})Z(\widetilde{Q}_{j})$
$=$ $Z(Q_{j-1})Z(Q_{j-2})Z(\widetilde{Q}_{j-3})U^{(2)}Z(\widetilde{Q}_{j-1})Z(\widetilde{Q}_{j})$
$=$ $\cdots$
$=$ $\left(\prod_{i\in I_{2}(j)}Z(Q_{i})\right)U^{(1)}\left(\prod_{i\in I_{1}(j)}Z(\widetilde{Q}_{i})\right),$
where the index sets $I_{1}(j)$ and $I_{2}(j)$ capture the specific pattern of the propagation, defined by
$I_{1}(j)\coloneqq\left\{\begin{aligned} &j-2k,\\ &j-2k-1\end{aligned}\;\middle|\;k\in\mathbb{N}_{0},j-2k\geq 1,j-2k-1\geq 1\right\},$ (S18)
and
$I_{2}(j)\coloneqq\left\{\begin{array}[]{l|l}j-2k-1,\,j-2k-2&\begin{array}[]{l}k\in\mathbb{N}_{0},\\ j-2k-1\geq 1,\\ j-2k-2\geq 1\end{array}\end{array}\right\}.$ (S19)
Thus, obtaining a measurement outcome $m_{j}=1$ on $\widetilde{Q}_{j}$ is equivalent to the ideal outcome $m_{j}=0$ up to the application of corrective $Z$ operations on the system qubits, specifically on those qubits whose indices belong to $I_{2}(j)$. The transfer matrix here has entries
$\mathcal{T}_{i,j}=\begin{cases}1,&\text{ if }i<j\text{ and }(j-2)\bmod 3\in\{1,2\},\\ 0,&\text{ otherwise.}\end{cases}$ (S20)
This prescription determines the corrective $Z$ operation on each system qubit $Q_{i}$ as a function of the parity of the measurement outcomes of a specific set of auxiliary systems.
A.0.2 Generalization and optimization
Deeper dynamic circuits—whether in 1D, on bipartite graphs, or on 2D grid topologies—can be analyzed using the formalism introduced earlier. For 1D topologies (see Fig. S1), fan-out ladders constructed with repeated $\mathrm{CX}$ layers (e.g., $\mathsf{FL}_{(2,3)}$, $\mathsf{FL}_{(2,3,4)}$, $\mathsf{FL}_{(1,2,4,5)}$, and $\mathsf{FL}_{(1,4,5,6)}$, where subscript indices indicate the relative distance between the target qubits from the control qubits within each fan-out gate) generally exhibit translational symmetry away from circuit boundaries despite partial cancellation of intermediate connections. For bipartite connectivity graphs (Fig. S2(a)), each system qubit interacts with multiple neighboring auxiliary qubits. Specifically, if system qubit $Q_{i}$ interacts with auxiliary qubits $\widetilde{Q}_{i-2}$, $\widetilde{Q}_{i-1}$, $\widetilde{Q}_{i}$, and $\widetilde{Q}_{i+1}$ (when available), the effective fan-out range increases more rapidly with additional layers. In 2D grid topologies, translational symmetry is typically lost. Fig. S3 shows illustrative examples of one-layer, two-layer, and square-grid topologies and their associated fan-out ladders. Our randomized dynamic circuits introduced earlier generalize these constructions, exhibiting an exponential increase in fan-out range with parameter $\mathfrak{r}_{2}$ when $\mathfrak{r}_{2}\ll\log(n)$. The precise set of target qubits is determined dynamically based on information flow within the circuit.
Feed-forward corrections across all these structures are consistently products of identity and Pauli-$Z$ operators, analyzed using the transfer matrix formalism. Corrections from the first fan-out ladder propagate through subsequent ladders via commutation relationships between Pauli-$Z$ and $\mathrm{CX}$ gates. After the final measurements, combined feed-forward corrections realize the complete fan-out staircase.
Now we discuss the synthesis and optimization of our randomized fan-out staircase circuits, specifically addressing whether our construction can be decomposed into shallower, more efficient circuits.
Figure S4: Comparison of $\mathrm{CX}$ ($\mathrm{CNOT}$) gate counts between original dynamic circuits, effective fan-out staircases after feed-forward corrections, and optimized effective circuits obtained via Gaussian elimination. (a) $\mathrm{CX}$ gate counts for fixed $\mathfrak{r}_{1}=1$ and varying $\mathfrak{r}_{2}$. (b) $\mathrm{CX}$ gate counts for fixed $\mathfrak{r}_{2}=1$ and varying $\mathfrak{r}_{1}$. Each data point represents the average over 50 randomly sampled instances, with error bars indicating standard deviations.
Finally, we compare the number of $\mathrm{CX}$ gates required by the original dynamic circuits, the corresponding effective fan-out staircases obtained via feed-forward corrections, and the optimized effective circuits derived through Gaussian elimination. We investigate whether these randomly generated fan-out staircases can be further optimized to reduce circuit complexity. Without loss of generality, numerical experiments are conducted with fixed parameters $n=300$ and $\mathfrak{D}=3$.
First, fixing $\mathfrak{r}_{1}=1$, we vary $\mathfrak{r}_{2}$, and sample 50 random circuit instances for each value. Results are shown in Fig. S4(a). Initially, for small $\mathfrak{r}_{2}$, the effective fan-out staircases require exponentially increasing numbers of $\mathrm{CX}$ gates compared to the original dynamic circuits. At larger values of $\mathfrak{r}_{2}$, effective circuits become significantly denser. However, beyond a certain threshold ($\mathfrak{r}_{2}>4$), applying Gaussian elimination to convert the circuits into optimized effective forms leads to reductions in $\mathrm{CX}$ count, revealing redundancy in the original staircase representation. Prior to this threshold, the Gaussian elimination method yields circuits with increased complexity, indicating no beneficial optimization.
Next, we fix $\mathfrak{r}_{2}=1$ and vary $\mathfrak{r}_{1}$ again sampling 50 random cases per parameter setting. As depicted in Fig. S4(b), the number of $\mathrm{CX}$ gates in logical fan-out staircases increases significantly more slowly.
Appendix B Appendix B: Measurement-driven IQP sampling
In the main text, we introduced measurement-driven IQP circuits, constructed by interleaving randomized measurement-enabled fan-out staircase circuits with layers of single-qubit Pauli-$Z$ rotations. This appendix provides relevant proofs of theoretical claims, complemented by additional technical discussions and illustrative examples.
We first prove Proposition 1, which formally states that Hamiltonian phase states with random binary architectures achieve anticoncentration for a linear number of measurement and rotation steps.
Proof of Proposition 1.
Consider the $n$-qubit Hamiltonian phase state vectors of the form
$\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle=\exp\left(\mathrm{i}\sum_{i=1}^{s}\vartheta_{i}\bigotimes_{j=1}^{n}Z^{\mathbf{A}_{i,j}}\right)\lvert+\rangle^{\otimes n},$ (S21)
where $\mathbf{A}\in\{0,1\}^{s\times n}$ is a binary matrix whose ith row specifies the qubits on which the diagonal gate $Z^{A_{i}}$ acts, and $\bm{\vartheta}=(\vartheta_{1},\dots,\vartheta_{s})$ with each $\vartheta_{i}\in[0,2\pi)$ drawn independently and uniformly. Here,
$|+\rangle^{\otimes n}\;\coloneqq\;\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle.$ (S22)
For a fixed architecture $\mathbf{A}$, its two-copy moment operator is defined as
$\rho^{(2)}_{\mathbf{A}}\;\coloneqq\;\mathbb{E}_{\bm{\vartheta}}\Bigl[\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle\langle\psi_{\mathbf{A},\bm{\vartheta}}\rvert^{\otimes 2}\Bigr].$ (S23)
Expressing $\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle$ in the computational basis, we obtain
$\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}\exp\Bigl(\mathrm{i}\sum_{i=1}^{s}\vartheta_{i}\,(-1)^{A_{i}\cdot x}\Bigr.)\,\lvert x\rangle,$ (S24)
where $(-1)^{A_{i}\cdot x}$ captures parity $(A_{i}\cdot x\mod 2)$. The second moment operator $|\psi_{\mathbf{A},\bm{\vartheta}}\rangle\langle\psi_{\mathbf{A},\bm{\vartheta}}|^{\otimes 2}$ then expands to
$\rho_{A,\bm{\vartheta}}^{\otimes 2}=$ $\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle\langle\psi_{\mathbf{A},\bm{\vartheta}}\rvert^{\otimes 2}$ (S25)
$=$ $\frac{1}{2^{2n}}\sum_{\begin{subarray}{c}x,x^{\prime}\\ y,y^{\prime}\end{subarray}}\exp\Bigl(\mathrm{i}\sum_{i=1}^{s}\vartheta_{i}\,\bigl[(-1)^{A_{i}\cdot x}+(-1)^{A_{i}\cdot x^{\prime}}{\\ }\hfil&-(-1)^{A_{i}\cdot y}-(-1)^{A_{i}\cdot y^{\prime}}\bigr]\Bigr.)\lvert x,x^{\prime}\rangle\langle y,y^{\prime}\rvert$.
Next, we average over the random phases $\vartheta_{i}\in[0,2\pi)$. Recall that
$\int_{0}^{2\pi}e^{\mathrm{i}\alpha\vartheta}\,\frac{d\vartheta}{2\pi}=\delta_{\alpha,0},$ (S26)
for integer $\alpha$. Thus, any term in the sum vanishes unless the exponent of each $\vartheta_{i}$ is effectively zero, i.e.,
$(-1)^{A_{i}\cdot x}\;+\;(-1)^{A_{i}\cdot x^{\prime}}\;=\;(-1)^{A_{i}\cdot y}\;+\;(-1)^{A_{i}\cdot y^{\prime}}\quad\forall\,i.$ (S27)
Equivalently, for each $i$, the row $A_{i}$ must fail to distinguish the pairs $(x,x^{\prime})$ and $(y,y^{\prime})$.
The identity and the swap contributions certainly survive. Indeed, if $(x,y)=(x^{\prime},y^{\prime})$ or $(x,y)=(y^{\prime},x^{\prime})$, the exponent vanishes and hence these terms survive with probability 1. Additional off-diagonal terms could survive only if all rows $A_{i}$ yield the same parity outcome for $(x,x^{\prime})$ and $(y,y^{\prime})$.
We now let each row $A_{i}$ of $\mathbf{A}$ be drawn independently and uniformly from $\{0,1\}^{n}$. Define the two-copy moment operator averaged over both $\mathbf{A}$ and $\bm{\vartheta}$:
$\rho^{(2)}\;\coloneqq\;\mathbb{E}_{{\mathbf{A}},{\bm{\vartheta}}}\Bigl[\lvert\psi_{\mathbf{A},\bm{\vartheta}}\rangle\langle\psi_{\mathbf{A},\bm{\vartheta}}\rvert^{\otimes 2}\Bigr].$ (S28)
We say that $\rho^{(2)}$ reproduces (or approximates) the second-moment operator of the uniform diagonal unitary distribution $\rho_{\mathrm{ideal}}^{(2)}$ if
$\|\rho^{(2)}-\rho_{\mathrm{ideal}}^{(2)}\|_{1}\;\leq\;\varepsilon.$ (S29)
If they match exactly, one obtains an exact diagonal 2-design Gross et al. (2007); if they are close in trace norm, one obtains an $\varepsilon$-approximate diagonal 2-design Nakata and Murao (2013).
For any fixed quadruple $(x,x^{\prime}),(y,y^{\prime})$ not in the identity or swap configuration, each random row $A_{i}$ has a non-vanishing probability $c<1$ of distinguishing these pairs via parity. That is, there is a constant probability that
$(-1)^{A_{i}\cdot x}+(-1)^{A_{i}\cdot x^{\prime}}\neq(-1)^{A_{i}\cdot y}+(-1)^{A_{i}\cdot y^{\prime}}.$ (S30)
Consequently, across s independent rows, the probability that all rows fail to distinguish them is at most $c^{s}$.
Since there are at most $2^{4n}$ such quadruples, a union bound implies that the expected number of surviving ”unwanted” off-diagonal contributions in the total operator is bounded by $2^{4n}\,\cdot\,c^{s}$. By choosing $s\gtrsim k\,n$ for some constant $k$, we make $2^{4n}\,c^{s}$ vanish exponentially in $n$. Hence, with high probability over choices of $\mathbf{A}$, no unwanted off-diagonal terms survive in $\rho_{\mathbf{A},\bm{\vartheta}}^{\otimes 2}$. In that case,
$\bigl\|\rho^{(2)}-\rho_{\mathrm{ideal}}^{(2)}\bigr\|_{1}\;\leq\;\varepsilon,$ (S31)
showing that the averaged ensemble indeed provides an $\varepsilon$-approximate diagonal 2-design when s is sufficiently large (linear in $n$).
Formally, let $\Upsilon(s)$ denote the joint distribution of random architectures $\mathbf{A}$ with s rows and random phases $\bm{\vartheta}$. Then the resulting ensemble
$\bigl\{|\psi_{\mathbf{A},\bm{\vartheta}}\rangle:(\mathbf{A},\bm{\vartheta})\sim\Upsilon(s)\bigr\},$ (S32)
forms an $\varepsilon$-approximate diagonal 2-design once $s\in\mathcal{O}(n)$. Since diagonal 2-designs are known to anti-concentrate Nakata et al. (2014); Hangleiter et al. (2018), it follows that the output distributions of these random Hamiltonian phase states exhibit the desired anti-concentration property Nakata et al. (2014). $\blacksquare$
Before proving Theorem 1, we state the following lemma which underpins the decomposition of multi-qubit diagonal interactions (i.e., multi-qubit $Z$-rotations) into sums of lower-body $Z$-interactions.
Lemma 1 (Circuit synthesis).
Let $\ell$ and $k$ be integers. Then any $r$-qubit diagonal rotation of the form
$\exp\left(\mathrm{i}\frac{\ell\pi}{2^{k}}Z_{1}Z_{2}\cdots Z_{r}\right)$ (S33)
can be synthesized by combining at most $k$-body diagonal (i.e., $Z$-type) interactions with appropriate phases, up to global phases that do not affect measurement outcomes.
Proof of Lemma 1.
Consider the operator
$U_{r}(\ell,k)=\exp\left(\mathrm{i}\frac{\ell\pi}{2^{k}}Z^{\otimes r}\right)$ (S34)
where $Z^{\otimes r}=Z_{1}\otimes Z_{2}\otimes\cdots\otimes Z_{r}$. Note that
$Z=I-2|1\rangle\langle 1|,$ (S35)
so
$Z^{\otimes r}=(I-2|1\rangle\langle 1|)^{\otimes r}.$ (S36)
We can express this expansion via binary strings $q\in\{0,1\}^{r}$. Specifically, let $(\frac{Z-I}{2})_{q}$ denote a tensor product of $r$ factors where each factor is $(\frac{Z-I}{2})$ if $q_{j}=1$ and $I$ otherwise (for position $j\in\{1,\dots,r\}$. Equivalently,
$(I-2|1\rangle\langle 1|)^{\otimes r}=\sum_{q\in\{0,1\}^{r}}(-2)^{|q|}\left(\frac{Z-I}{2}\right)_{q},$ (S37)
where $|q|$ is the Hamming weight of $q$. Thus,
$U_{r}(\ell,k)=\exp\left(\mathrm{i}\frac{\ell\pi}{2^{k}}\sum_{q\in\{0,1\}^{r}}(-2)^{|q|}\left(\frac{Z-I}{2}\right)_{q}\right).$ (S38)
To confine ourselves to at most $k$-body terms, we observe that each string $q$ of weight $|q|>k$ contributes a rotation angle that is an integer multiple of $2\pi$ (since $\ell\pi/2^{k}$ is the base angle and any multiple of $2\pi$ is a global phase that cancels out in quantum mechanics). Consequently, we may omit all terms corresponding to $|q|>k$.
Rewriting each $(\frac{Z-I}{2})_{q}$ as a tensor product of $I$ and $Z$, we see that each surviving term of Hamming weight $|q|\leq k$ is effectively a $|q|$-body $Z$-interaction. In total, the exponent can be broken down into a linear combination of diagonal operators of body-size at most $k$. Hence, $U_{r}(\ell,k)$ can be decomposed into a product of multi-qubit diagonal operators, each of which involves at most $k$ qubits, up to an irrelevant global phase.
$\blacksquare$
Remark 1 (Alternative circuit synthesis strategies).
A similar decomposition strategy, leveraging expansions in terms of lower-weight $Z$-type interactions, has been discussed in Refs. Shepherd (2010); Rajakumar et al. .
To illustrate the above construction, let us consider a concrete example for $r$-qubit rotations of the form
$\exp\left(\mathrm{i}\frac{\ell\pi}{8}Z_{1}Z_{2}\cdots Z_{r}\right),$ (S39)
with $\ell\in\left\{0,1,\ldots,2^{k}-1\right\}$. One finds that it can be realized by composing into the following steps.
  1. 1.
    Single-qubit $Z$-rotations:
    $\exp\left(\mathrm{i}\alpha Z_{i}\right)\quad\text{ for each }i\in\{1,2,\ldots,r\}$ (S40)
    with
    $\alpha=\frac{(r^{2}-5r+6)\ell\pi}{4}.$ (S41)
  2. 2.
    Two-qubit $Z$-rotations,
    $\exp\left(\mathrm{i}\beta Z_{i}Z_{j}\right)\quad\text{ for all distinct }i,j,$ (S42)
    with
    $\beta=\frac{(3-r)\ell\pi}{8}.$ (S43)
  3. 3.
    Three-qubit $Z$ rotations
    $\exp\left(\mathrm{i}\gamma Z_{i}Z_{j}Z_{k}\right)\quad\text{ for all distinct }i,j,k$ (S44)
    with
    $\gamma=\frac{\ell\pi}{8}.$ (S45)
Details of the exact angle assignments follow from analyzing the expansion of $(Z-I)^{\otimes r}$. (or, equivalently, from the binomial expansion viewpoint in Lemma 1) and ensuring terms of weights 1, 2, and 3 combine to yield the correct net rotation on each subset of qubits.
Proof of Theorem 1.
The measurement-driven IQP circuits $\mathcal{C}_{\mathrm{MD}}$ interleave fan-out staircases with parallel single-qubit diagonal gates $\exp(\mathrm{i}\vartheta_{i,j}Z_{j})$ for preparing Hamiltonian phase states. Any effective multi-qubit diagonal rotation in $\mathcal{C}_{\mathrm{MD}}$ ultimately reduces to a product of multi-qubit diagonal unitaries of the form
$\exp\left(\mathrm{i}\frac{\ell\pi}{2^{k}}Z_{j_{1}}Z_{j_{2}}\cdots Z_{j_{r}}\right)$ (S46)
acting on distinct qubit subsets, for some integers $\ell$ and $r$.
From Lemma 1, each such $r$-qubit rotation with angle $\ell\pi/2^{k}$ is equivalent (up to a global phase) to combining at most $k$-body diagonal $Z$-interactions with integer multiples of $2\pi$. Hence, when all angles $\bm{\vartheta}$ lie in $\{\ell\,\pi/2^{k}\}$, $\mathcal{C}_{\mathrm{MD}}$ can be fully synthesized by commuting $Z$-interactions of at most $k$ qubits each. Because all gates remain diagonal in the computational basis, they commute pairwise. The net effect is equivalent to a $k$-local IQP with commuting $Z$-gates of weight at most $k$.
Therefore, the entire protocol, though it may appear to be a constant-depth dynamic circuit with intermediate measurements, is unitarily equivalent to a polynomial-depth standard IQP circuit containing multi-qubit diagonal gates. By ”pulling forward” all feed-forward corrections and combining them with the single-qubit or multi-qubit rotations, we see that the final action is local with up to $k$-body terms.
$\blacksquare$
In the main text, we introduced Criterion 1 to rigorously characterize the statistical randomness of a binary architecture matrix $\mathbf{A}\in\mathbb{Z}_{2}^{s\times n}$. Here, we elaborate in detail on the three underlying conditions comprising this criterion:
Figure S5: Comparison of statistical randomness metrics under 2D lattice connectivity. Panels (a–c) depict eigenvalue distributions of standardized covariance matrices from random $\mathrm{CX}$ circuits without auxiliary qubits at fixed depth 96 for system sizes $n=14\times 14,28\times 28,42\times 42$. Panels (d–f) show eigenvalue distributions from measurement-driven randomized fan-out staircases with auxiliary qubits at fixed depth 18 (parameters: $\mathfrak{D}=3$, $\mathfrak{r}_{1}=2$, $\mathfrak{r}_{2}=1$) for system sizes $n=20\times 20,28\times 28,42\times 42$. Panels (g–i) illustrate distributions of pairwise row Hamming distances for random $\mathrm{CX}$ circuits without auxiliary qubits at depth 96, and panels (j–l) show corresponding distributions for measurement-driven circuits at depth 18, where half the qubits serve as auxiliary qubits.
Spectral (eigenvalue) requirement.
The standardized covariance matrix
$Y_{\mathbf{A}}=\frac{1}{s}\bigl(2\mathbf{A}-1\bigr)\bigl(2\mathbf{A}-1\bigr)^{\top}$ (S47)
should exhibit an empirical eigenvalue distribution closely matching the Marchenko–Pastur (MP) law Marchenko and Pastur (1967). Formally, for an aspect ratio $\gamma=n/s$ (or vice versa, depending on which dimension grows faster), the MP distribution is given by the probability density function
$\rho_{\mathrm{MP}}^{(\gamma)}(\lambda)\;=$ $\max\Bigl\{0,1-\tfrac{1}{\gamma}\Bigr\}\,\delta(\lambda)$ (S48)
$+\;\frac{1}{2\pi\,\gamma\,\lambda}\,\sqrt{\,\bigl(\lambda-a_{-}\bigr)\,\bigl(a_{+}-\lambda\bigr)\,}$
$\quad\text{for}\;\lambda\in[\,a_{-},\,a_{+}]\,$,
where
$a_{\pm}\;=\;\bigl(1\,\pm\,\sqrt{\gamma}\bigr)^{2}.$ (S49)
In the limit $s,n\to\infty$ with fixed ratio $\gamma$, random matrices of appropriate form converge almost surely to this distribution for their eigenvalue spectra.
Fig. S5(a–f) illustrates spectral density comparisons at fixed circuit depths across increasing system sizes, contrasting matrices generated from measurement-driven fan-out staircases with auxiliary qubits against those from local random $\mathrm{CX}$ circuits without auxiliary qubits. As the system size grows at a fixed depth, matrices from random $\mathrm{CX}$ circuits increasingly deviate from the ideal MP distribution, whereas eigenvalues from measurement-driven circuits remain closely aligned.
Hamming (combinatorial) requirement:
The distributions of Hamming weights—defined as the number of ones within each row $\mathbf{A}_{i,:}$ or each column $\mathbf{A}_{:,j}$—should closely follow binomial distributions $\mathrm{Binomial}(n,p)$ or $\mathrm{Binomial}(s,p)$, respectively, where $p\approx 1/2$. This implies that the entries of each row or column should behave approximately as independent random bits with equal probability.
Moreover, the pairwise Hamming distances between distinct rows or columns should similarly approximate a binomial distribution with parameter $p=1/2$. Explicitly, the pairwise Hamming distance between two distinct rows $\mathbf{A}_{i,:}$ and $\mathbf{A}_{j,:}$ is
$d_{\mathrm{H}}(\mathbf{A}_{i,:},\;\mathbf{A}_{j,:})\;=\;\sum_{\ell=1}^{n}\bigl|\mathbf{A}_{i,\ell}-\mathbf{A}_{j,\ell}\bigr|,$ (S50)
and should concentrate near $n/2$. Analogous conditions apply for columns. Satisfying these conditions ensures no systematic bias or correlations occur within the matrix entries.
Fig. S5(g–l) demonstrates illustrative distributions of row–row Hamming distances across various system sizes. As system size increases at fixed depth, the random $\mathrm{CX}$ circuits without auxiliary qubits progressively deviate from the theoretical binomial distribution, while the measurement-driven circuits consistently maintain alignment with the theoretical expectation.
Rank requirement over $\mathrm{GF}(2)$:
A fundamental question in random matrix theory over finite fields is: What is the probability that an $n\times n$ binary matrix, whose entries are i.i.d. with probability $1/2$ of being 0 or 1, has full rank nkn-k? Ref. Kolchin (1999) provides an explicit formula for this probability in the case of $\mathrm{GF}(2)$. Concretely, if we consider an $n\times n$ binary matrix $\mathbf{A}$ and let
[…]
then
$P_{n,k}=2^{-k^{2}}$ $\biggl(\prod_{\ell=0}^{n-k-1}\bigl(1-2^{-(n-\ell)}\bigr)\biggr)$ (S52)
$\cdot\biggl(\sum_{0\leq i_{1}\leq\cdots\leq i_{k}\leq n-k}2^{-(i_{1}+\cdots+i_{k})}\biggr).$
Here, $k$ is a fixed nonnegative integer representing the “rank deficit” from $n$. In other words, $k=0$ corresponds to full rank, $k=1$ corresponds to rank n1n-1, etc.
When $k$ remains fixed and $n\to\infty$, $P_{n,k}$ converges to a limiting value
$\widetilde{P}_{k}$ $\;=\;\lim_{n\to\infty}P_{n,k}$ (S53)
= $\;=2^{-\,k^{2}}\biggl(\,\prod_{i\geq k+1}\bigl(1-2^{-\,i}\bigr)\biggr)\,\biggl(\,\prod_{\,i=1}^{\,k}\bigl(1-2^{-\,i}\bigr)^{-1}\biggr).$
Numerically,
$\widetilde{P}_{0}\;\approx 028879,\;\;\widetilde{P}_{1}\;\approx 057758,\;\;\widetilde{P}_{2}\;\approx 012835,$ (S54)
and their summation is approximately 0.99471.
Within our setting, we require that submatrices of $\mathbf{A}$ retain high rank over $\mathrm{GF}(2)$ with large probability in order to certify algebraic randomness. Since certain sampled matrix sizes $n$ may not be large enough to apply the strict asymptotics, we relax the criterion to demand that at least 90% of randomly chosen submatrices achieve a rank within two of the maximum possible. Fig. S6 compares these probabilities for architecture matrices $\mathbf{A}$ generated by (i) random $\mathrm{CX}$ circuits without auxiliary qubits, and (ii) measurement-driven fan-out staircases with auxiliary qubits. In particular, we plot $\widetilde{P}_{0}+\widetilde{P}_{1}+\widetilde{P}_{2}$ against system size for both methods, revealing that measurement-driven matrices maintain near-full rank across a broader range of sizes and depths.
Figure S6: $\widetilde{P}_{0}+\widetilde{P}_{1}+\widetilde{P}_{2}$ (the probability that a random submatrix of the $n\times n$ binary matrix has rank over $\mathrm{GF}(2)$ at least n2n-2) as a function of system size for architecture matrices $\mathbf{A}$ produced by (i) local random $\mathrm{CX}$ circuits without auxiliary qubits (red) and (ii) measurement-driven fan-out staircases with auxiliary qubits (blue).
Appendix C Appendix C: Geometrically local quantum circuit cost in two-dimensional lattices
In the main text, we generalized the concept of quantum circuit cost and its entanglement-based lower bound, introduced in Ref. Eisert (2021), to two-dimensional (2D) lattice architectures. Quantum circuit cost quantifies the minimal resources required to implement a unitary operation by integrating the entangling capabilities of geometrically local gates. Here, we provide rigorous definitions and explicitly derive a lower bound for this cost in terms of the sum of bipartite entanglement entropies.
Consider a 2D lattice of qubits initially prepared in a product state $|0\rangle^{\otimes n}$. Suppose a target state vector $|\psi\rangle=U|0\rangle^{\otimes n}$ is generated by a unitary $U$ composed exclusively of nearest-neighbor two-qubit interactions. Following Ref. Eisert (2021), the 2D geometrically local circuit cost $C^{(2D)}_{g}(U)$ is defined as the minimal total integrated interaction strength required to realize $U$. Formally, let
$U=\mathcal{P}\exp\left(-\mathrm{i}\int_{0}^{1}\mathcal{H}(s)\,ds\right)$ (S55)
with a time-dependent Hamiltonian
$\mathcal{H}(s)=\sum_{e\in\mathcal{E}}y_{e}(s)\,O_{e},$ (S56)
where each $O_{e}$ is a normalized two-qubit operator acting across lattice edge $e$, the circuit cost is then
$C_{g}^{(2D)}(U):=\inf_{{y_{e}(s)}}\int_{0}^{1}\sum_{e\in\mathcal{E}}|y_{e}(s)|ds.$ (S57)
To derive the entanglement-based lower bound, consider bipartitions of the lattice defined by vertical cuts $\ell_{x}^{(a)}$ between columns and horizontal cuts $\ell_{y}^{(b)}$ between rows, each partitioning the lattice into two distinct regions. The entanglement entropy for a bipartition $\ell$ is defined as
[…]
where $\rho_{L(\ell)}$ is the reduced density matrix on one side of the cut.
The fundamental principle underlying the lower bound is the small incremental entangling principle Mariën et al. (2016); Eisert (2021), which limits the rate of entanglement generation by local operations. Specifically, at any time $s$, the growth rate of entanglement entropy across a cut $\ell$ is bounded as
$\frac{d}{ds}S(\psi(s),\ell)\leq\eta\sum_{e\in\mathcal{E}:\,O_{e}\text{ crosses }\ell}|y_{e}(s)|$ (S59)
with $\eta$ an $\mathcal{O}(1)$ constant dependent solely on geometry and operator norms.
Integrating this inequality over the full evolution yields
$S(\psi,\ell)\leq\eta\int_{0}^{1}\sum_{e\in\mathcal{E}:\,O_{e}\text{ crosses }\ell}|y_{e}(s)|\,ds.$ (S60)
Since each two-qubit gate lies exactly on one nearest-neighbor cut—either horizontal or vertical—its entangling contribution is counted precisely once in either the horizontal or vertical entropy sums. Therefore, summation over all vertical and horizontal cuts gives
$C_{g}^{(2D)}(U)\geq\frac{1}{\eta}\left(\sum_{\ell_{x}}S(\psi,\ell_{x})+\sum_{\ell_{y}}S(\psi,\ell_{y})\right).$ (S61)
We define the entanglement-based quantity (scaled by a constant $1/\eta$)
$\xi(\psi):=\frac{1}{\eta}\left(\sum_{\ell_{x}}S(\psi,\ell_{x})+\sum_{\ell_{y}}S(\psi,\ell_{y})\right),$ (S62)
leading directly to the inequality
$\xi(\psi)\leq C_{g}^{(2D)}(U).$ (S63)
Thus, the sum of bipartite entropies across all lattice cuts, appropriately scaled, constitutes a rigorous lower bound for the geometrically local quantum circuit cost in 2D lattice systems, generalizing the one-dimensional result from Ref. Eisert (2021).
Appendix D Appendix D: Measurement‑based multibody XY quantum reservoir
For a fixed measurement-enabled architecture matrix $\mathbf{A}\in\{0,1\}^{s\times n}$ (here $s=\Theta(n)$), define the Pauli strings
$P_{A_{i}}^{(\mu)}\;:=\;\bigotimes_{j=1}^{n}\sigma^{\mu\,\mathbf{A}_{i,j}}_{j},\,\mu\in\{x,y\},\;i=1,\dots,s.$ (S64)
One Floquet step of the reservoir is the ordered product
$U_{\mathsf{MD}}(\tau)=\prod_{\mu=x,y}\exp\!\left(\mathrm{i}\,\tau\sum_{i=1}^{s}c^{\mu}_{i}\,P_{A_{i}}^{(\mu)}\right).$ (S65)
After $K=T/(2\tau)$ identical steps the measurement‑driven reservoir channel reads
$\mathcal{R}_{\mathrm{MD}}^{(T)}(\rho)=\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{K}\,\rho\,\bigl(U_{\mathsf{MD}}(\tau)\bigr)^{\dagger K}.$ (S66)
The measurement‑enabled XY reservoir possesses vastly higher expressivity than any geometrically local Hamiltonian and, in principle, can extract many kinds of global information. For Theorem 2, however, it suffices to exhibit one concrete instance with only one Floquet cycle where this advantage produces a constant‑size read‑out gap.
Proof of Theorem 2. Let $G=(V,E)$ be a connected graph with maximum degree $\deg(G)=\mathcal{O}(1)\leq\Delta$. It contains a simple path of length at least $n/\Delta$. By selecting appropriate vertices along this path, pick a triplet
$\mathcal{S}=\{i,j,k\}\subset V$ (S67)
such that the pairwise graph distances satisfy
$\operatorname{dist}(i,j),\,\operatorname{dist}(j,k),\,\operatorname{dist}(k,i)\geq\frac{n}{3\Delta}.$ (S68)
Furthermore, choose a measurement-driven architecture matrix $\mathbf{A}\in\{0,1\}^{s\times n}$ so that there exists a row $r$ with
$\exists\,r:\;\mathbf{A}_{r,i}=\mathbf{A}_{r,j}=\mathbf{A}_{r,k}=1.$ (S69)
Let
$H_{\mathcal{S}}:=H_{i}H_{j}H_{k},\qquad X_{\mathcal{S}}:=X_{i}X_{j}X_{k},$ (S70)
and set the observable $O:=Z_{i}Z_{j}Z_{k}$. For $\ell\in\{0,1\}$, define
[…]
Then
[…]
Now we choose the xx‑sector couplings as $\widetilde{c}_{r}^{x}:=\pi/4$ and, for each row $i\neq r$, we set the new coefficients
$\tilde{c}_{i}^{x}:=\begin{cases}\frac{\varepsilon}{n}c_{i}^{x}&|\mathbf{A}_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{x}&|\mathbf{A}_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2),\end{cases}$ (S73)
with $\varepsilon\ll 1$, $c_{i}^{x}$ being the initially sampled coefficients, $|\mathbf{A}_{i}|_{\mathcal{S}}$ is the overlap of row $A_{i}$ with $\mathcal{S}$.
For each row $i$, we choose the $y$-sector couplings
$\tilde{c}_{i}^{y}:=\begin{cases}\frac{\varepsilon}{n}\,c_{i}^{y}&|\mathbf{A}_{i}|_{\mathcal{S}}=1\,(\mathrm{mod}\,2),\\ c_{i}^{y}&|\mathbf{A}_{i}|_{\mathcal{S}}=0\,(\mathrm{mod}\,2).\end{cases}$ (S74)
Rows with an odd overlap $|\mathbf{A}_{i}|_{\mathcal{S}}=1,3$ give Pauli strings that anticommute with the read‑out observable $O=Z_{i}Z_{j}Z_{k}$; their couplings are suppressed by the factor $\varepsilon/n$ with $0<\varepsilon\ll 1$. For any such string $\mathbf{A}_{i}$ we have the first‑order estimate
$\exp\Bigl(\frac{\mathrm{i}\varepsilon}{n}P^{(\mu)}_{\mathbf{A}_{i}}\Bigr.)O\exp\Bigl(-\frac{\mathrm{i}\varepsilon}{n}P^{(\mu)}_{\mathbf{A}_{i}}\Bigr.)=O+\mathcal{O}\Bigl(\frac{\varepsilon}{n}\Bigr),$ (S75)
so their contribution to the expectation of $O$ is negligible. Rows with an even overlap $|\mathbf{A}_{i}|_{\mathcal{S}}=0,2$ commute with $O$ and retain their sampled strength. With these choices one Floquet period leads to
[…]
Thus a single measurement‑based Floquet cycle already gives the required constant read‑out gap.
For any nearest‑neighbor Hamiltonian generating a unitary $V$ over the same time $T=\mathcal{O}(1)$, the Lieb–Robinson bound gives
$V^{\dagger}OV=\tilde{O}_{R}+\tilde{E},\quad\|\tilde{E}\|=\mathcal{O}(2^{-n}),$ (S77)
[…]
Equations (S76) and (S78) establish the constant–vs.– exponential separation claimed in Theorem 2. $\blacksquare$
The same construction applies to other sector choices: for an XZ (resp. YZ) reservoir pick $O=Y_{i}Y_{j}Y_{k}$ (resp. $O=X_{i}X_{j}X_{k}$), and choose the heavy row in a sector that anticommutes with $O$. For the full XYZ reservoir, any choice of heavy sector with $O$ taken in the complementary Pauli suffices. The odd/even‑overlap suppression and the Lieb–Robinson contrast carry over verbatim, yielding the same constant read‑out gap.
Appendix E Appendix E: Additional numerical benchmarks for the measurement‑based quantum reservoir
This section presents a series of comparative studies on the measurement-driven quantum reservoir computing (QRC) model, analyzing its performance under different conditions and its robustness to measurement errors. In every comparison, we hold the data-generation pipeline, feature budget, and evolution-time budget constant, varying only a single factor at a time to isolate its effect.
Figure S7: Performance and robustness of the measurement-based QRC. (a, b) Heavy-hex and square-lattice layouts used to define the reservoir architectures. (c) Classifier-independent performance: The measurement-based reservoir (filled markers) consistently outperforms the local Ising baseline (open markers) for both ridge regression (blue) and kk‑nearest neighbors (red) readouts. (d) Pauli sector comparison: The measurement-based reservoir achieves high accuracy regardless of the chosen Pauli interaction sector $S\in\{\mathrm{XY},\mathrm{XZ},\mathrm{YZ},\mathrm{XYZ}\}$ using the SVM readout. (e) Robustness to measurement error: The reservoir’s accuracy shows only a graceful degradation with increasing per-qubit measurement error, demonstrating feasibility with near-term hardware. Each point is an average over 20 independent architectures (10 for each layout).
We generate input states from the extended SSH chain on $n=16$ system qubits, which exhibits three distinct phases (trivial, topological, and symmetry-broken). For each phase, we uniformly sample an eigenstate from the lowest 20 energy levels. Each sampled state is then prepared for the reservoir by applying a weak, local coherent perturbation: for every qubit $i$, we independently draw two angles
$\theta_{i}^{(x)}\sim\mathcal{N}(0,0.03),\quad\theta_{i}^{(z)}\sim\mathcal{N}(0,0.03),$ (S79)
and apply the rotation $R_{x}\big(\theta_{i}^{(x)}\big)\,R_{z}\big(\theta_{i}^{(z)}\big)$.
The reservoir dynamics are governed by Hamiltonians with $s=16$ terms, where all coupling strengths are sampled i.i.d. from $\mathcal{N}(0,1/2)$. The system undergoes ten identical Floquet cycles. We consider two hardware-motivated layouts: heavy-hex and square lattices [Figs. S7(a,b)]. For each layout, we sample 10 independent architecture matrices (either randomized fan-out staircases or degree-matched local edge sets), meaning each data point represents an average over 20 distinct architectures. The corresponding standard deviations are shown as error bars.
The output features are the set of single-site expectation values,
$\mathbf{f}(T)=\big(\langle\sigma_{1}^{z}\rangle,\ldots,\langle\sigma_{n}^{z}\rangle\big),$ (S80)
which are measured at the end of each Floquet cycle up to a total evolution time $T$. Each expectation value is estimated from 8192 shots. Unless specified otherwise, our simulation includes a symmetric per-qubit readout flip error of $\varepsilon_{\mathrm{meas}}=5\times 10^{-3}$. The final dataset is class-balanced, and we use an identical train/test split across all methods. Features are standardized using statistics from the training set only.
We benchmark two families of reservoirs. The first, our local baseline, is a transverse-field Ising (TFI) model on the given hardware layout,
$\mathcal{H}_{\mathrm{TFI}}=\sum_{\langle i,j\rangle}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{x}_{i}.$ (S81)
The second is the measurement-based multibody reservoir, implemented with mid-circuit measurements and Pauli-frame feed-forward as described in the main text. This method generates global interactions in a chosen Pauli sector $S\in\{\mathrm{XY},\mathrm{XZ},\mathrm{YZ},\mathrm{XYZ}\}$ from an architecture matrix $\mathbf{A}\in\{0,1\}^{s\times n}$:
$U_{\mathsf{MD}}(\tau)=\prod_{\mu\in S}\exp\left(\mathrm{i}\,\tau\sum_{i=1}^{s}c_{i}^{\mu}\bigotimes_{j=1}^{n}\sigma_{j}^{\mu\,\mathbf{A}_{i,j}}\right).$ (S82)
For a fair comparison, all benchmarks match the total evolution time $T$ and the number of Hamiltonian terms.
To assess the classifier-independence of our results, we supplement the SVM analysis with two distinct models: ridge regression (a linear model) and $k$-nearest neighbors ($k=21$), a non-parametric method. Figure S7(c) shows that the measurement-based reservoir consistently achieves higher accuracy than the local TFI baseline, regardless of the chosen classifier. The baseline’s performance, in contrast, shows little improvement with increasing evolution time $T$. Since all other experimental conditions are identical, this strongly indicates that the observed advantage stems from the superior quantum feature map generated by the measurement-driven dynamics—specifically its global mixing and input-dependent nonlinearity—and is not an artifact of the classical readout method.
Figure S7(d) tests the performance dependence on the specific Pauli sector used for the multibody interactions. We compare reservoirs generated from four different sectors $(S=\mathrm{XY},\mathrm{XZ},\mathrm{YZ},\mathrm{XYZ})$, keeping all other parameters fixed. The results show that all four choices yield quantitatively similar learning curves, achieving high classification accuracy at moderate evolution times. This indicates that the performance gain is robust and not tied to a specific interaction basis, but rather stems from the general ability to create global, multibody gates via measurement and feed-forward, as formalized in our expressivity separation (Theorem 2).
We also investigate the reservoir’s robustness to measurement noise, a key factor for near-term implementation. As shown in Fig. S7(e), we test the classification accuracy against a per-qubit readout error rate increasing from 0.5% to 2.5%. The performance exhibits a graceful degradation. For both longer ($T=1$) and shorter ($T=0.01$) evolution times, the drop in accuracy is minor and remains well within the statistical uncertainty from architecture sampling. This resilience to noise at levels typical of current quantum hardware underscores the practical feasibility of our approach.
In summary, all benchmarks presented here were performed under strictly controlled conditions: the input data, feature space $\{\langle\sigma_{i}^{z}\rangle\}$, shot budget, evolution time, and Hamiltonian parameter counts were held constant for each comparison. These controlled tests, combined with the analysis in the main text showing a significant performance drop when feed-forward is disabled, provide strong evidence for our central claim. They collectively demonstrate that the measurement-and-feed-forward mechanism itself—which enables the creation of global and nonlinear feature maps—is the primary driver of the observed performance advantage.